OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14 (3 points) No solutions x = -1, x = 8.3 x = 0, x = 7 x = -1, x = 7

OpenStudy (raffle_snaffle):

solve for x

OpenStudy (ipwnbunnies):

^^

OpenStudy (ipwnbunnies):

Great help. 8)

OpenStudy (anonymous):

3|x - 3| = 12->|x - 3| = 4 a) -|x - 3| = 4 -> -x = 4 - 3 = 1 -> x = ... b) |x - 3| = 4 -> x = 3 + 4 = ....

OpenStudy (eg145341):

It's been a while since I've dealt with this kind of stuff but I think you solve for x by doing this... 3| x - 3| + 2 = 14 - 2 -2 3|x - 3| = 14 Multiply 3 by x and -3 3x - 9 = 14 + 9 +9 3x = 23 /3 /3 \[x \approx 7\]

OpenStudy (anonymous):

So C